OpenStudy (loser66):

\(\int \dfrac{sin^2x}{7e^x}dx\) Please, help

OpenStudy (loser66):

@zepdrix

ganeshie8 (ganeshie8):

use the identity \(\sin^2x = \frac{1-\cos (2x)}{2}\), the integral becomes \[\frac{1}{14}\int e^{-x}\, dx-\frac{1}{14}\int e^{-x}\cos(2x)\, dx\]

OpenStudy (loser66):

Yes, I did

ganeshie8 (ganeshie8):

There are several ways to evaluate an integral of form \(\int e^{ax}\cos(bx)\, dx\) there is really a very neat method if you're okay with complex numbers

OpenStudy (loser66):

ah, you want to express it in term of Euler?

OpenStudy (loser66):

I am ok with any method, please. show me. I can't get the answer as what wolfram does.

OpenStudy (loser66):

ok, go ahead, please

ganeshie8 (ganeshie8):

Yes : \[\large e^{-x}\cos(2x) = \mathcal{R} (e^{-x+i2x})\]

ganeshie8 (ganeshie8):

we're done, integrating \(e^{-x+i2x}\) is a piece of cake

ganeshie8 (ganeshie8):

\[\begin{align} \int e^{-x}\cos(2x) \,dx &= \mathcal{R} \int e^{-x+i2x}\,dx\\~\\ &= \mathcal{R}~ \dfrac{e^{-x+i2x}}{-1+2i}\\~\\ \end{align}\] just get the real part of that expression and yeah don't forget the integration constant..

OpenStudy (loser66):

Got you, thank you. Much appreciate. :)

ganeshie8 (ganeshie8):

np :)